Suppose that X ~ χ2(m), Y-2(n), and X and Y are independent. Is Y-X ~ χ2 if
Answer #1

The MGF of Xis My (t) = (1-2)-2 and Y ~ χ。4 The MGF of Yis , My() = (1-2) Let Z-Y-x The MGF of Z is = (1-2)-3(1+2)T which isn